OpenStudy (anonymous):

Solve the system by elimination. x+y+z=7 x+2y+2z=11 3x+y-z=9 x=? , y=? , z=?

OpenStudy (anonymous):

add equation one to equation three if you choose

OpenStudy (anonymous):

that leaves me with 4x+2y=16 and x+2y+2z=11 correct?

OpenStudy (anonymous):

yea, now you can solve 4x+2y=16 for either x or y

OpenStudy (anonymous):

y=-2x+8

OpenStudy (anonymous):

x=3,y=2,z=2

OpenStudy (anonymous):

2y=16-4x wouldn't turn into y=8-2x?

OpenStudy (anonymous):

ha ha.. sorry, yea, ...it does.

OpenStudy (anonymous):

oh woosh... I was worried I was way off base lol

OpenStudy (anonymous):

thanks Ryan

OpenStudy (anonymous):

why not thank me

OpenStudy (anonymous):

and thank you love math ;)

OpenStudy (anonymous):

you kind of snuck in here

OpenStudy (anonymous):

you got it from here?